# Conditional Denoising Diffusion Probabilistic Models for Data Reconstruction Enhancement in Wireless Communications

### v2

@@ Abstract
In this paper, conditional denoising diffusion probabilistic models (DDPMs) are proposed to enhance the data transmission and reconstruction over wireless channels. The underlying mechanism of DDPM is to decompose the data generation process over the so-called “denoising” steps. Inspired by this, the key idea is to leverage the generative prior of diffusion models in learning a “noisy-to-clean” transformation of the information signal to help enhance data reconstruction. The proposed scheme could be beneficial for communication scenarios in which a prior knowledge of the information content is available, e.g., in multimedia transmission. Hence, instead of employing complicated channel codes that reduce the information rate, one can exploit diffusion priors for reliable data reconstruction, especially under extreme channel conditions due to low signal-to-noise ratio (SNR), or hardware-impaired communications. The proposed DDPM-assisted receiver is tailored for the scenario of wireless image transmission using MNIST dataset. Our numerical results highlight the reconstruction performance of our scheme compared to the conventional digital communication, as well as the deep neural network (DNN)-based benchmark. It is also shown that more than 10 dB improvement in the reconstruction could be achieved in low SNR regimes, without the need to reduce the information rate for error correction.
@@
AI-native wireless, diffusion priors, generative AI, network resilience, wireless AI, conditional denoising diffusion models.
@@ I-A Motivation & Background
With the emergence of deep generative models, the realm of artificial intelligence (AI) has witnessed a paradigm shift towards the novel concept of generative AI that can facilitate the development of AI-based systems [2]. At the same time, from data communication and networking perspective, “connected intelligence” is envisioned as the most significant driving force in the sixth generation (6G) of communications—AI and machine learning (AI/ML) algorithms are envisioned to be widely incorporated into 6G wireless networks, realizing “AI-native” communication systems [3, 5, 4, 6]. This underscores the need for novel AI/ML-based solutions to be tailored to meet the requirements of emerging communication scenarios.
The evolution of diffusion models [7], as the new paradigm in deep generative models, can be regarded as one of the main factors contributing to the recent breakthroughs in generative AI, that have already showcased notable success with some of the popular examples such as DALL.E 2 by OpenAI, [^1] https://openai.com/dall-e-2 and ImageGen by Google Brain. [^2] https://imagen.research.google/ Diffusion models have realized unprecedented results in different applications, such as computer vision, natural language processing (NLP), data analysis and synthesis, and medical imaging [8]. Please see [9] and references therein for a comprehensive survey on diffusion models. The core concept of diffusion modelling is that if we could develop an ML model that is capable of learning the systematic decay of information, due to noise/distortions, then it should be possible to “reverse” the process and recover the information back from the noisy/erroneous data. This is fundamentally different from conventional generative models, like generative adversarial networks (GANs) and variational autoencoders (VAEs). Accordingly, the close underlying relation between the key concepts on how diffusion models work and the problems in wireless communication systems has motivated us to carry out this research.
@@ I-B Related Works
Ongoing research on diffusion models encompasses both theoretical advancements and practical applications across different domains of computer science [9]. However, there have been only a few papers in wireless communication literature that have started looking into the potential merits of diffusion-based generative models for wireless systems [10, 11, 12, 13, 14, 15]. The authors in [10] study a workflow for utilizing diffusion models in wireless network management. They study the incentive mechanisms for generating contracts using diffusion models in mobile AI-generated content services. They exploit the exploration capability of diffusion models in Q-learning algorithms for network management. Denoising diffusion model is utilized in [11] to generate synthetic channel realizations conditioned on the message information. The authors tackle the problem of differentiable channel model within the gradient-based training process of end-to-end ML-based communications. The results highlight the performance of diffusion models as an alternative to generative adversarial network (GAN)-based schemes. It is shown in [11] that GANs experience unstable training and less diversity in generation performance, due to their adversarial training nature. However, diffusion models maintain a more stable training process and a better generalization during inference. Noise-conditioned score network (NCSN)-based channel estimation is studied in [12] and [13] for multi-input-multi-output (MIMO) wireless communication scenario. The authors estimate the gradient of the high-dimensional log-prior of wireless channels, using RefineNet neural network architecture. Posterior sampling method is further applied to make the generated channel estimations consistent with the pilot signals measured at the receiver. Diffusion models have also been incorporated into the so-called semantic communication systems in [15, 14]. More specifically, deep learning-based joint source-channel coding (Deep-JSCC) is combined with diffusion models to complement semantic data transmission schemes with a generative component. The initial results show that the perceptual quality of reconstruction can be improved in such scenarios. Nevertheless, those schemes in [15, 14] propose to employ two different types of deep learning models each maintaining a distinct objective function (while their optimization also depends on each other), i.e., the autoencoder for semantic transmission and reception, and the diffusion models for reconstruction enhancement. This can impose serious concerns in terms of the computational overhead to the network. In addition, the output signal of a neural encoder (i.e., the channel input) does not necessarily follow the standard format of modulation signals, which makes it challenging to employ semantic transmission ideas in a real-world communication scenario.
@@ I-C Our Contributions
In this paper, we propose conditional denoising diffusion probabilistic models (DDPMs) for enhancing the wireless image transmission in digital communication schemes. Different from GANs and VAEs, the underlying mechanism of DDPM is to first transform the ground-truth signal into an isotropic Gaussian distribution in the forward diffusion process, and then try to regenerate data over the so-called “denoising” steps by removing the noise introduced in the diffusion process. DDPMs in their vanilla format are based upon the assumption that the denoising process starts with isotropic Gaussian noise [7]. However, in a practical wireless system, the received/decoded signals are a degraded version of the ground-truth information signals, not necessarily a pure isotropic Gaussian noise, while the degradation is different in different distortion conditions such as different SNRs and transceiver specifications. To address this, we employ conditional diffusion probabilistic models, in which the noisy degraded signals are incorporated into the denoising framework of the diffusion model. This way, the diffusion model learns to reconstruct the ground-truth information from the degraded received signal, instead of a pure noise.
@@
In our scheme, the generative prior of conditioned DDPMs is exploited to enhance data reconstruction at the receiver of a communication system. This is particularly pronounced in extreme channel conditions with poor connectivity due to low signal-to-noise ratio (SNR), or hardware-impaired communications. In such scenarios, error correcting codes might not be able to correct the mismatches, or complicated codes might be needed at the cost of decreasing the information rate to correct the errors. Instead, the proposed diffusion-aided scheme can be utilized for reliable data reconstruction without the need for decreasing the information rate. This can be tailored for the communication scenarios, in which a prior knowledge of the information signals is available that can be exploited by a DDPM to learn the “denoising” process.
@@
The proposed DDPM-assisted receiver is employed in conjunction with the NVIDIA Sionna simulator [20], by plugging the trained diffusion model to the receiver side, highlighting the compatibility of the scheme with a practical communication system. Moreover, in contrast to [15, 14], the proposed scheme does not rely on any additional autoencoder model. Numerical results highlight that the resilience of the wireless system can be improved against practical non-idealities such as hardware impairments and channel distortions. We compare the reconstruction performance of our scheme to the conventional digital communications without diffusions, and also the deep neural network (DNN)-based benchmark [18]. It is also shown that more than 10 dB improvement in the reconstruction could be achieved in low SNR regimes, without the need to reduce the information rate for error correction.
@@ I-D Paper Organization and Notations
The rest of the paper is organized as follows. In Section II, we first introduce the concept of DDPMs, the forward, and the reverse diffusion processes, together with the main formulas and the corresponding loss functions. We also try to point out insightful intuitions behind the main formulas of the DDPM framework. Our system model is introduced in Section III. We first provide the generic formulation of a communication system under non-linear HWIs. We then address the proposed conditional DDPM by modifying the vanilla diffusion model. Numerical experiments are provided in Section IV. Finally, Section V concludes the paper.
@@
II Preliminaries on Denoising Diffusion Probabilistic Models
Diffusion models are a new family of state-of-the-art probabilistic generative models which are inspired by non-equilibrium thermodynamics [16]. They are comprised of two processes, namely the forward diffusion process and the parametric reverse diffusion process, as illustrated in Fig. 1. During the forward process, samples are diffused using Gaussian kernels to add noise until the fully-distorted signals follow an isotropic Gaussian distribution. Then in the reverse process, the model tries to “decode” the data via “denoising” the perturbed samples in a hierarchical fashion. In this paper, we focus on DDPMs a the new state-of-the-art family of diffusion models. In what follows, we study the concept of DDPMs together with the main formulas and the corresponding loss functions. We also try to point out insightful intuitions behind the underlying formulas of the DDPM framework.
II-A Forward Diffusion Process
Let ${\bm{x}}_{0}$ be a data sample from some distribution ${q}({\bm{x}}_{0})$, i.e., ${\bm{x}}_{0}\sim{q}({\bm{x}}_{0})$. For a finite number, $T$, of time-steps, the forward diffusion process $q({\bm{x}}_{t}|{\bm{x}}_{t-1})$ is defined by adding Gaussian noise at each time-step $t\in[T]$ according to a known “variance schedule” $10<\beta_{1}<\beta_{2}<\cdots<\beta_{T}<$ 1. This can be formulated as
@@ II-A Forward Diffusion Process
Invoking (2), the data sample gradually loses its distinguishable features as the time-step goes on, where with $T\!\rightarrow\!\infty$, ${\bm{x}}_{T}$ approaches an isotropic Gaussian distribution with covariance matrix ${\bf\Sigma}\!=\!\sigma^{2}\bm{I}$ for some $\sigma\!>\!0$ [7]. According to (1), each new sample at time-step $t$ can be drawn from a conditional Gaussian distribution with mean vector ${\bm{\mu}}_{t}=\sqrt{1-\beta_{t}}\bm{x}_{t-1}$ and covariance matrix ${\bf\Sigma}^{2}_{t}=\beta_{t}\bf I$. Hence, the forward process is realized by sampling from a Gaussian noise $\bm{\epsilon}_{t-1}\sim{\cal N}(\bf 0,I)$ and setting
@@
A useful property for the forward process in (3) is that we can sample ${\bm{x}}_{t}$ at any arbitrary time-step $t$ in a closed-form expression, through recursively applying the reparameterization trick from ML literature [19], and considering the sum properties of two Gaussian random variables. This results in the following formulation
@@
$\bar{\alpha}_{t}\!$ $=\!\prod_{i=1}^{t}(1-\alpha_{i}),\quad\alpha_{t}=1-\beta_{t}.$ (6)
Notably, (4) implies that one can directly sample ${\bm{x}}_{t}$, $t\in[T]$ from $x_{0}$.
II-B Parametric Reverse Diffusion Process
Now the problem is to reverse the process in (4) and sample from $q(\bm{x}_{t-1}|\bm{x}_{t})$, so that we regenerate the true samples from some Gaussian noise $\bm{x}_{T}$. According to [16], for $\beta_{t}$ small enough, $q(\bm{x}_{t-1}|\bm{x}_{t}),\forall t\in[T]$ also follows Gaussian distribution. However, we cannot easily estimate the distribution, since it requires knowing the distribution of all possible data samples (or equivalently exploiting the entire dataset) to be able to calculate the conditional probability. Hence, to approximate the conditional probabilities and run the reverse diffusion process, we need to learn a probabilistic model $p_{\bm{\theta}}(\bm{x}_{t-1}|\bm{x}_{t})$ that is parameterized by ${\bm{\theta}}$. Accordingly, the following expressions can be written
$p_{\bm{\theta}}(\bm{x}_{t-1}|\bm{x}_{t})$ $\sim\mathcal{N}(\bm{x}_{t-1};\bm{\mu}_{\bm{\theta}}(\bm{x}_{t},t),\bm{\Sigma}_{\bm{\theta}}(\bm{x}_{t},t)),$ (7)
$p_{\bm{\theta}}(\bm{x}_{0:T})$ $=p(\bm{x}_{T})\prod^{T}_{t=1}p_{\bm{\theta}}(\bm{x}_{t-1}|\bm{x}_{t}).$ (8)
As can be seen from (7), the mean and variance of the reverse process are also conditioned on the time-step $t$. Now the problem simplifies to learning the mean vector $\bm{\mu}_{\bm{\theta}}(x_{t},t)$ and the covariance matrix $\bm{\Sigma}_{\bm{\theta}}(x_{t},t)$ for the probabilistic model $p_{\bm{\theta}}(\cdot)$, where a neural network can be trained to approximate (learn) the reverse process with ${\bm{\theta}}$ showing the parameters of the neural network.
Before proceeding with the details of learning the reverse diffusion process, we note that if we condition the reverse process on ${\bm{x}}_{0}$, this conditional probability becomes tractable. The intuition behind this could be explained as follows. A painter (our generative model) requires a reference image ${\bm{x}}_{0}$ to be able to gradually draw a picture. Hence, when we have ${\bm{x}}_{0}$ as a reference, we can take a small step backwards from noise to generate the data samples, and the reverse step (drawing a picture in our metaphor) would be formulated as $q({\bm{x}}_{t-1}|{\bm{x}}_{t},{\bm{x}}_{0})$. Mathematically speaking, by knowing the conditional probabilities of $q({\bm{x}}_{t}|{\bm{x}}_{0})$ and $q({\bm{x}}_{t-1}|{\bm{x}}_{0})$, and utilizing Bayes rule, we can derive $q({\bm{x}}_{t-1}|{\bm{x}}_{t},{\bm{x}}_{0})$ with a similar expression to (7)

### v3

@@ Abstract
In this paper, conditional denoising diffusion probabilistic models (CDiffs) are proposed to enhance the data transmission and reconstruction over wireless channels. The underlying mechanism of diffusion models is to decompose the data generation process over the so-called “denoising” steps. Inspired by this, the key idea is to leverage the generative prior of diffusion models in learning a “noisy-to-clean” transformation of the information signal to help enhance data reconstruction. The proposed scheme could be beneficial for communication scenarios in which a prior knowledge of the information content is available, e.g., in multimedia transmission. Hence, instead of employing complicated channel codes that reduce the information rate, one can exploit diffusion priors for reliable data reconstruction, especially under extreme channel conditions due to low signal-to-noise ratio (SNR), or hardware-impaired communications. The proposed CDiff-assisted receiver is tailored for the scenario of wireless image transmission using MNIST dataset. Our numerical results highlight the reconstruction performance of our scheme compared to the conventional digital communication, as well as the deep neural network (DNN)-based benchmark. It is also shown that more than 10 dB improvement in the reconstruction could be achieved in low SNR regimes, without the need to reduce the information rate for error correction.
@@
AI-native wireless, diffusion priors, generative AI, wireless AI, conditional denoising diffusion models, reliable communication.
Figure 1: Data pipeline of our experimental communication setup for CDiff-based image reconstruction.
@@ I-A Motivation & Background
With the emergence of deep generative models, the realm of artificial intelligence (AI) has witnessed a paradigm shift towards the novel concept of generative AI that can facilitate the development of AI-based systems [2, 3, 4]. At the same time, from data communication and networking perspective, “connected intelligence” is envisioned as the most significant driving force in the sixth generation (6G) of communications—AI and machine learning (AI/ML) algorithms are envisioned to be widely incorporated into 6G wireless networks, realizing “AI-native” communication systems [5, 7, 6, 8]. This underscores the need for novel AI/ML-based solutions to be tailored to meet the requirements of emerging communication scenarios.
The evolution of diffusion models [9], as the new paradigm in deep generative models, can be regarded as one of the main factors contributing to the recent breakthroughs in generative AI, that have already showcased notable success with some of the popular examples such as DALL.E 2 by OpenAI, [^1] https://openai.com/dall-e-2 and ImageGen by Google Brain. [^2] https://imagen.research.google/ Diffusion models have realized unprecedented results in different applications, such as computer vision, natural language processing (NLP), data analysis and synthesis, and medical imaging [10]. Please see [11] and references therein for a comprehensive survey on diffusion models. The core concept of diffusion modeling is that if we could develop an ML model that is capable of learning the systematic decay of information, due to noise/distortions, then it should be possible to “reverse” the process and recover the information back from the noisy/erroneous data. This is fundamentally different from conventional generative models, like generative adversarial networks (GANs) and variational autoencoders (VAEs). Accordingly, the close underlying relation between the key concepts on how diffusion models work and the problems in wireless communication systems has motivated us to carry out this research.
@@ I-B Related Works
Ongoing research on diffusion models encompasses both theoretical advancements and practical applications across different domains of computer science [11]. However, there have been only a few papers in wireless communication literature that have started looking into the potential merits of diffusion-based generative models for wireless systems [12, 13, 4, 3, 14, 15, 16, 17]. The authors in [12] study a workflow for utilizing diffusion models in wireless network management. They study the incentive mechanisms for generating contracts using diffusion models in mobile AI-generated content services. They exploit the exploration capability of diffusion models in Q-learning algorithms for network management. Denoising diffusion model is utilized in [13] to generate synthetic channel realizations conditioned on the message information. The authors tackle the problem of differentiable channel model within the gradient-based training process of end-to-end ML-based communications. The results highlight the performance of diffusion models as an alternative to GAN-based schemes. It is shown in [13] that GANs experience unstable training and less diversity in generation performance, due to their adversarial training nature. However, diffusion models maintain a more stable training process and a better generalization during inference. Noise-conditioned score network (NCSN)-based channel estimation is studied in [14] and [15] for multi-input-multi-output (MIMO) wireless communication scenario. The authors estimate the gradient of the high-dimensional log-prior of wireless channels, using RefineNet neural network architecture. Posterior sampling method is further applied to make the generated channel estimations consistent with the pilot signals measured at the receiver. Diffusion models have also been incorporated into the so-called semantic communication systems in [17, 16]. More specifically, deep learning-based joint source-channel coding (Deep-JSCC) is combined with diffusion models to complement semantic data transmission schemes with a generative component. The initial results show that the perceptual quality of reconstruction can be improved in such scenarios. Nevertheless, those schemes in [17, 16] propose to employ two different types of deep learning models each maintaining a distinct objective function (while their optimization also depends on each other), i.e., the autoencoder for semantic transmission and reception, and the diffusion models for reconstruction enhancement. This can impose serious concerns in terms of the computational overhead to the network. In addition, the output signal of a neural encoder (i.e., the channel input) does not necessarily follow the standard format of modulation signals, which makes it challenging to employ semantic transmission ideas in a real-world communication scenario.
@@ I-C Our Contributions
In this paper, we propose conditional denoising diffusion probabilistic models (DDPM) for enhancing the wireless image transmission in digital communication schemes. Different from GANs and VAEs, the underlying mechanism of diffusion models is to first transform the ground-truth signal into an isotropic Gaussian distribution in the forward diffusion process, and then try to regenerate data over the so-called “denoising” steps by removing the noise introduced in the diffusion process. DDPMs in their vanilla format are based upon the assumption that the denoising process starts with isotropic Gaussian noise [9]. However, in a practical wireless system, the received/decoded signals are a degraded version of the ground-truth information signals, not necessarily a pure isotropic Gaussian noise, while the degradation is different in different distortion conditions such as different SNRs and transceiver specifications. To address this, we propose conditional diffusion models (CDiff), in which the noisy degraded signals are incorporated into the denoising framework of the diffusion model. This way, the diffusion model learns to reconstruct the ground-truth information from the degraded received signal, instead of a pure noise.
@@
In our scheme, the generative prior of conditioned DDPMs is exploited to enhance data reconstruction at the receiver of a communication system. This is particularly pronounced in extreme channel conditions with poor connectivity due to low signal-to-noise ratio (SNR), or hardware-impaired communications. In such scenarios, error correcting codes might not be able to correct the mismatches, or complicated codes might be needed at the cost of decreasing the information rate to correct the errors. Instead, the proposed diffusion-aided scheme can be utilized for reliable data reconstruction without the need for decreasing the information rate. This can be tailored for the communication scenarios, in which a prior knowledge of the information signals is available that can be exploited by a DDPM to learn the “denoising” process. Our scheme is not simply a direct application of the seminal papers in diffusion model research. We do not directly apply the seminal DDPM paper [9] into a communication system. Rather, different from [9], we incorporate the wireless “conditions” into the diffusion model via employing a wireless CDiff.
@@
The proposed DDPM-assisted receiver is employed in conjunction with the NVIDIA Sionna simulator [^3] https://github.com/NVlabs/sionna, by plugging the trained diffusion model to the receiver side. Data communication pipeline of our scheme is shown in Fig. 1, which highlights the compatibility of the scheme with a practical communication system. Moreover, in contrast to [17, 16], the proposed scheme does not rely on any additional autoencoder model. Numerical results highlight that the robustness of the wireless system can be improved against practical non-idealities such as hardware impairments and channel distortions. We compare the reconstruction performance of our scheme to the conventional digital communications without diffusions, and also the deep neural network (DNN)-based benchmark [18]. It is also shown that more than 10 dB improvement in the reconstruction could be achieved in low SNR regimes, without the need to reduce the information rate for error correction.
@@ I-D Paper Organization and Notations
The rest of the paper is organized as follows. In Section II, we first introduce the concept of DDPMs, the forward, and the reverse diffusion processes, together with the main formulas and the corresponding loss functions. We also try to point out insightful intuitions behind the main formulas of the DDPM framework. Our system model is introduced in Section III. We first provide the generic formulation of a communication system under RF impairments. We then address the proposed conditional DDPM by modifying the vanilla diffusion model. Numerical experiments are provided in Section IV. Finally, Section V concludes the paper.
@@
II Investigation of Denoising Diffusion Probabilistic Models
In this section, we provide insights into the underlying mechanisms of the DDPM model. We also provide a comprehensive overview of the theoretical foundations and practical implications of the DDPM model, offering a deeper understanding of its potential impact on communication systems.
@@ II Investigation of Denoising Diffusion Probabilistic Models
II-A Theoretical Foundations
Diffusion models are comprised of two processes, namely the forward diffusion process and the parametric reverse diffusion process, as illustrated in Fig. 2. During the forward process, samples are diffused using Gaussian kernels to add noise until the fully-distorted signals follow an isotropic Gaussian distribution. Then in the reverse process, the model tries to “decode” the data via “denoising” the perturbed samples in a hierarchical fashion.
II-A1 Forward Diffusion Process
Let ${\bm{x}}_{0}$ be a data sample from some distribution ${q}({\bm{x}}_{0})$, i.e., ${\bm{x}}_{0}\sim{q}({\bm{x}}_{0})$. For a finite number, $T$, of time-steps, the forward diffusion process $q({\bm{x}}_{t}|{\bm{x}}_{t-1})$ is defined by adding Gaussian noise at each time-step $t\in[T]$ according to a known “variance schedule” $10<\beta_{1}<\cdots<\beta_{T}<$ 1. This is formulated as
@@ II-A1 Forward Diffusion Process
Invoking (2), the data sample gradually loses its distinguishable features as the time-step goes on, where with $T\!\rightarrow\!\infty$, ${\bm{x}}_{T}$ approaches an isotropic Gaussian distribution. According to (1), each new sample at time-step $t$ can be drawn from a conditional Gaussian distribution with mean vector ${\bm{\mu}}_{t}=\sqrt{1-\beta_{t}}\bm{x}_{t-1}$ and covariance matrix ${\bf\Sigma}^{2}_{t}=\beta_{t}\bf I$. Hence, the forward process is realized by sampling from a Gaussian noise $\bm{\epsilon}_{t-1}\sim{\cal N}(\bf 0,I)$ and setting
@@
By recursively applying the reparameterization trick from ML literature [20], we can sample ${\bm{x}}_{t}$ at any arbitrary time-step $t$ in a closed-form expression. This results in
@@
where $\bar{\alpha}_{t}=\!\prod_{i=1}^{t}(1-\alpha_{i})$ and $\alpha_{t}=1-\beta_{t}$. Note that (4) implies that we can directly sample ${\bm{x}}_{t}$ from $x_{0}$ for all $t\in[T]$.
II-A2 Parametric Reverse Diffusion Process
Now the problem is to reverse the process in (4) and sample from $q(\bm{x}_{t-1}|\bm{x}_{t})$, so that we regenerate the true samples from some Gaussian noise $\bm{x}_{T}$. However, we cannot easily estimate the distribution, since it requires knowing the distribution of all possible data samples (or equivalently exploiting the entire dataset). Hence, to approximate the conditional probabilities and run the reverse diffusion process, we need to learn a probabilistic model $p_{\bm{\theta}}(\bm{x}_{t-1}|\bm{x}_{t})$ that is parameterized by ${\bm{\theta}}$. Accordingly, the following expressions can be written
$p_{\bm{\theta}}(\bm{x}_{t-1}|\bm{x}_{t})$ $\sim\mathcal{N}(\bm{x}_{t-1};\bm{\mu}_{\bm{\theta}}(\bm{x}_{t},t),\bm{\Sigma}_{\bm{\theta}}(\bm{x}_{t},t)).$ (6)
Now the problem simplifies to learning the mean vector $\bm{\mu}_{\bm{\theta}}(x_{t},t)$ and the covariance matrix $\bm{\Sigma}_{\bm{\theta}}(x_{t},t)$ for the probabilistic model $p_{\bm{\theta}}(\cdot)$, where a neural network (parameterized by ${\bm{\theta}}$) can be trained to approximate (learn) the reverse process. When we have ${\bm{x}}_{0}$ as a reference, we can take a small step backwards from noise to generate the data samples, and the reverse step would be formulated as $q({\bm{x}}_{t-1}|{\bm{x}}_{t},{\bm{x}}_{0})$. Mathematically speaking, by knowing the conditional probabilities of $q({\bm{x}}_{t}|{\bm{x}}_{0})$ and $q({\bm{x}}_{t-1}|{\bm{x}}_{0})$, and utilizing Bayes rule, we can derive $q({\bm{x}}_{t-1}|{\bm{x}}_{t},{\bm{x}}_{0})$ with a similar expression to (6):